# Replica Thermodynamic Trade-off Relations: Entropic Bounds on Network Diffusion and Trajectory Observables

### v1

@@ Methods
First, we consider the single-process case. The probability of the trajectory $\Gamma$ [Eq. (4)] is given by [61]
$\mathcal{P}\left(\Gamma\right)=p_{\mu}(t=0)\mathcal{P}\left(\Gamma\mid\mu_{0}\right),$ (31)
@@
where we used Eqs. (37) and (36). Equation (39) is Eq. (12). Taking $s\to\infty$ and $r=1$, we have
$\mathbb{E}\left[|N^{(\otimes K)}|^{s}\right]^{r/(s-r)}=\max_{\Gamma^{(\otimes K)}}|N^{(\otimes K)}(\Gamma^{(\otimes K)})|.$ (40)
Using Eq. (40) in Eq. (39), we obtain Eq. (14) in the main text.
.2 Derivation of Eqs. (16) and (17)
@@
$\Lambda^{(\otimes 2)}(\Gamma^{\otimes 2})\equiv\begin{cases}0&N^{(1)}(\Gamma^{(1)})=N^{(2)}(\Gamma^{(2)})\\ 1&\mathrm{otherwise}\end{cases}.$ (41)
$\Lambda^{(\otimes 2)}$ is interpreted as follows. Consider the trajectories $\Gamma^{(1)}$ and $\Gamma^{(2)}$ of two independent processes, and take the observable $N^{(k)}(\Gamma^{(k)})$, defined in Eq. (11), on these two processes. Then $\Lambda^{(\otimes 2)}$ returns 0 only when the values of the observable $N$ in the two processes agree, $N^{(1)}(\Gamma^{(1)})=N^{(2)}(\Gamma^{(2)})$. When there is no jump in the first and second processes, we trivially obtain $N^{(1)}(\Gamma_{\varnothing}^{(1)})=N^{(2)}(\Gamma_{\varnothing}^{(2)})=0$. Then the observable $\Lambda^{(\otimes 2)}$ satisfies
${\Lambda}^{(\otimes 2)}(\Gamma_{\varnothing}^{\otimes 2})=0,$ (42)
which is the condition of Eq. (11). Let $\mathcal{Z}$ be the set of all possible values that the observable $N(\Gamma)$ can take. Then, the expectation becomes the following:
$\mathbb{E}\left[\Lambda^{(\otimes 2)}(\Gamma^{\otimes 2})\right]$
$=1-\sum_{z\in\mathcal{Z}}P(N^{(1)}(\Gamma^{(1)})=z)P(N^{(2)}(\Gamma^{(2)})=z)$
$=1-\sum_{z\in\mathcal{Z}}P(N(\Gamma)=z)^{2}.$ (43)
Moreover, since ${\Lambda}^{(\otimes 2)}$ returns either $1$ or 0, we have
$[{\Lambda}^{(\otimes 2)}(\Gamma^{\otimes 2})]_{\max}=1.$ (44)
Substituting Eqs. (43) and (44) into Eq. (14), we obtain
$\sum_{z\in\mathcal{Z}}P(N(\Gamma)=z)^{2}\geq e^{-2\tau\mathfrak{a}(t=0)}.$ (45)
The derivation above was presented for the case $K=2$. It can be straightforwardly extended to any integer $K\geq 2$, yielding
$\sum_{z\in\mathcal{Z}}P(N(\Gamma)=z)^{K}\geq e^{-K\tau\mathfrak{a}(t=0)},$ (46)
@@
$\Xi^{(\otimes 2)}(\Gamma^{\otimes 2})\equiv\begin{cases}0&X^{(1)}(\tau)=X^{(2)}(\tau)\\ 1&\mathrm{otherwise}\end{cases}.$ (47)
Here, $X^{(k)}(\tau)$ is the state of the $k$-th process at time $t=\tau$. Therefore, Eq. (47) is 0 if the final state of the first and the second processes agree. In this study, when considering a random walker on a network, we assume that the initial state of the random walk is $B_{\mu_{0}}$. This implies that, when there is no jump within the interval $0,\tau]$, the final state at time $t=\tau$ is $X^{(1)}(\tau)=X^{(2)}(\tau)=B_{\mu_{0}}$. Therefore, the observable $\Xi^{(\otimes 2)}$ satisfies
$\Xi^{(\otimes 2)}(\Gamma_{\varnothing}^{(\otimes 2)})=0,$ (48)
which satisfies the condition given in Eq. (11). The expectation of $\Xi^{(\otimes 2)}$ is given by
$\mathbb{E}\left[\Xi^{(\otimes 2)}(\Gamma^{(\otimes 2)})\right]$
$=1-\sum_{\mu}P(X^{(1)}(\tau)=B_{\mu}^{(1)})P(X^{(2)}(\tau)=B_{\mu}^{(2)})$
$=1-\sum_{\mu}p_{\mu}(\tau)^{2}.$ (49)
Again, the maximum of $\Xi^{(\otimes 2)}$ is given by
$[\Xi^{(\otimes 2)}(\Gamma^{\otimes 2})]_{\max}=1.$ (50)
Substituting Eqs. (49) and (50) into Eq. (14), we obtain
$\sum_{\mu}p_{\mu}(\tau)^{2}\geq e^{-2\tau\mathfrak{a}(t=0)}.$ (51)
The derivation above was presented for the case $K=2$. It can be straightforwardly extended to any integer $K\geq 2$, yielding
$\sum_{\mu}p_{\mu}(\tau)^{K}\geq e^{-K\tau\mathfrak{a}(t=0)},$ (52)
@@
.3 Derivation of Eq. (28)
We show the derivation of Eq. (28), which is the quantum generalization of Eq. (12). Similarly to the classical case, let $\mathfrak{p}_{\mathrm{q}}(\tau)$ be the probability that there is no jump within the interval $0,\tau]$ in $\Gamma_{\mathrm{q}}$. From the result in Ref. [45], the following relation holds:
$\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2}\leq\mathfrak{p}_{\mathrm{q}}(\tau),$ (53)
@@
We consider the $K$ replica scenario, where there are $K$ identical processes. Let $\mathfrak{p}^{(\otimes K)}_{\mathrm{q}}(\tau)$ denote the probability that there is no jump in the $K$ replica processes. Since each process is independent, from Eq. (53), $\mathfrak{p}^{(\otimes K)}_{\mathrm{q}}(\tau)$ satisfies
$\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K}\leq\mathfrak{p}_{\mathrm{q}}^{(\otimes K)}(\tau).$ (55)
If no jump occurs in the interval $0,\tau]$, then the observable satisfies $N_{\mathrm{q}}^{(\otimes K)}(\Gamma_{\mathrm{q}}^{(\otimes K)})=0$ by the condition in Eq. (26). The converse, however, does not necessarily hold. Therefore, the probability that $N_{\mathrm{q}}^{(\otimes K)}(\Gamma_{\mathrm{q}}^{(\otimes K)})=0$ is bounded from below by
$P\left[N_{\mathrm{q}}^{(\otimes K)}(\Gamma_{\mathrm{q}}^{(\otimes K)})=0\right]\geq\mathfrak{p}_{\mathrm{q}}^{(\otimes K)}(\tau).$ (56)
@@
$\frac{\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}{\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}$
$\leq P(|N_{\mathrm{q}}^{(\otimes K)}|>0)$
$=1-P(|N_{\mathrm{q}}^{(\otimes K)}|=0)$
$\leq 1-\mathfrak{p}_{\mathrm{q}}^{(\otimes K)}(\tau)$
@@
where we used Eqs. (55) and (56). Equation (57) proves Eq. (28) in the main text. Similarly to Eq. (40), taking $s\to\infty$ and $r=1$, we have
$\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|^{s}\right]^{r/(s-r)}=\max_{\Gamma_{\mathrm{q}}^{(\otimes K)}}|N_{\mathrm{q}}^{(\otimes K)}(\Gamma_{\mathrm{q}}^{(\otimes K)})|.$ (58)
By using Eq. (58) in Eq. (57), we reproduce Eq. (29).
.4 Derivation of Eq. (30)
We show the derivation of Eq. (30), which directly parallels the derivation of Eq.(17). Specifically, let us consider the following observable:
$\Lambda_{\mathrm{q}}^{(\otimes 2)}(\Gamma_{\mathrm{q}}^{\otimes 2})\equiv\begin{cases}0&N_{\mathrm{q}}^{(1)}(\Gamma_{\mathrm{q}}^{(1)})=N_{\mathrm{q}}^{(2)}(\Gamma_{\mathrm{q}}^{(2)})\\ 1&\mathrm{otherwise}\end{cases}.$ (59)
Then $\Lambda_{\mathrm{q}}^{(\otimes 2)}$ returns 0 only when the values of the observable $N_{\mathrm{q}}$ in the two processes agree, $N_{\mathrm{q}}^{(1)}(\Gamma_{\mathrm{q}}^{(1)})=N_{\mathrm{q}}^{(2)}(\Gamma_{\mathrm{q}}^{(2)})$. Then the observable $\Lambda_{\mathrm{q}}^{(\otimes 2)}$ satisfies
$\Lambda_{\mathrm{q}}^{(\otimes 2)}(\Gamma_{\mathrm{q}\varnothing}^{\otimes 2})=0,$ (60)
which is the condition of Eq. (26). Let $\mathcal{Z}_{\mathrm{q}}$ be the set of all possible values that the observable $N_{\mathrm{q}}(\Gamma_{\mathrm{q}})$ can take. Then, the expectation becomes the following:
$\mathbb{E}\left[\Lambda_{\mathrm{q}}^{(\otimes 2)}(\Gamma_{\mathrm{q}}^{\otimes 2})\right]$
$=1-\sum_{z\in\mathcal{Z}_{\mathrm{q}}}P(N_{\mathrm{q}}^{(1)}(\Gamma_{\mathrm{q}}^{(1)})=z)P(N_{\mathrm{q}}^{(2)}(\Gamma_{\mathrm{q}}^{(2)})=z)$
$=1-\sum_{z\in\mathcal{Z}_{\mathrm{q}}}P(N_{\mathrm{q}}(\Gamma_{\mathrm{q}})=z)^{2}.$ (61)
Moreover, since ${\Lambda}^{(\otimes 2)}$ returns either $1$ or 0, we have
$[\Lambda_{\mathrm{q}}^{(\otimes 2)}(\Gamma_{\mathrm{q}}^{\otimes 2})]_{\max}=1.$ (62)
By considering a general $K$ case and substituting Eqs. (61) and (62) into Eq. (29), we obtain
$\sum_{z\in\mathcal{Z}_{\mathrm{q}}}P(N_{\mathrm{q}}(\Gamma_{\mathrm{q}})=z)^{K}\geq\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K},$ (63)
@@
.5 Quantum dynamical activity
We briefly review the quantum dynamical activity, a quantum generalization of the dynamical activity defined in Eq. (13).
@@
$\mathcal{A}(\tau)\equiv\int_{0}^{\tau}\sum_{m}\mathrm{Tr}\left[L_{m}\rho(t)L_{m}^{\dagger}\right]dt.$ (64)
In the classical limit, this expression reduces to the classical dynamical activity:
$\mathcal{A}_{\mathrm{cl}}(\tau)=\int_{0}^{\tau}\mathfrak{a}(t)dt=\int_{0}^{\tau}\sum_{\mu^{\prime},\mu(\mu\neq\mu^{\prime})}W_{\mu^{\prime}\mu}p_{\mu}(t)dt.$ (65)
@@
$\mathcal{B}(\tau)=\mathcal{A}(\tau)+\mathcal{C}(\tau),$ (66)
@@
Here, $\check{H}(t)\equiv e^{\mathcal{L}^{\dagger}t}H$ is the Hamiltonian $H$ in the Heisenberg picture, and $\mathcal{L}^{\dagger}$ is the adjoint Liouvillian,

### v2

@@ IV Conclusion
Suppose that the initial state in the trajectory $\Gamma$ is $B_{\mu_{0}}$. The probability of the trajectory $\Gamma$ [Eq. (4)] is given by [67]
$\mathcal{P}\left(\Gamma\right)=p_{\mu_{0}}(t=0)\mathcal{P}\left(\Gamma\mid\mu_{0}\right),$ (73)
@@
where we used Eqs. (78) and (79). Equation (81) is Eq. (17).
Appendix C Derivation of Eqs. (27) and (34)
@@
$\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})\equiv\begin{cases}0&N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma^{(K)})\\ 1&\mathrm{otherwise}\end{cases}$ (82)
$\overline{\Lambda}^{(\otimes K)}_{\circ}$ is similar to Eq. (58). However, the cases in which Eqs. (58) and (82) return 0 and 1 are reversed. $\overline{\Lambda}_{\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{\circ}$ in $K$ processes agree, $N_{\circ}^{(1)}(\Gamma^{(1)})=N_{\circ}^{(2)}(\Gamma^{(2)})=\cdots N_{\circ}^{(K)}(\Gamma^{(K)})$. When there is no jump in all $K$ processes, we trivially obtain $N_{\circ}^{(1)}(\Gamma_{\varnothing}^{(1)})=N_{\circ}^{(2)}(\Gamma_{\varnothing}^{(2)})=\cdots=N_{\circ}^{(K)}(\Gamma_{\varnothing}^{(K)})=0$. Then the observable $\overline{\Lambda}_{\circ}^{(\otimes K)}$ satisfies
$\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})=0,$ (83)
which is the condition of Eq. (13). Let $\mathcal{Z}_{\circ}$ be the set of all possible values that the observable $N_{\circ}(\Gamma)$ can take. Then, the expectation becomes the following:
$\mathbb{E}\left[\overline{\Lambda}_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})\right]$ $=1-\sum_{z\in\mathcal{Z}_{\circ}}\prod_{k=1}^{K}P(N_{\circ}^{(k)}(\Gamma^{(k)})=z)$
$=1-\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}(\Gamma)=z)^{K}.$ (84)
Substituting Eq. (84) into Eq. (17), we obtain
$\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}(\Gamma)=z)^{K}\geq e^{-K\tau\mathfrak{a}(t=0)}.$ (85)
@@
$\overline{\Xi}^{(\otimes K)}(\Gamma^{(\otimes K)})\equiv\begin{cases}0&X^{(1)}(\tau)=X^{(2)}(\tau)=\cdots=X^{(K)}(\tau)\\ 1&\mathrm{otherwise}\end{cases}.$ (86)
Here, $X^{(k)}(\tau)$ is the state of the $k$-th process at time $t=\tau$. Therefore, Eq. (86) is 0 if the final states of the $K$ processes agree. In this study, when considering a random walker on a network, we assume that the initial state of the random walk is $B_{\mu_{0}}$. This implies that, when there is no jump within the interval $0,\tau]$, the final state at time $t=\tau$ is $X^{(1)}(\tau)=X^{(2)}(\tau)=\cdots=X^{(K)}(\tau)=B_{\mu_{0}}$. Therefore, the observable $\overline{\Xi}^{(\otimes K)}$ satisfies
$\overline{\Xi}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})=0,$ (87)
which satisfies the condition given in Eq. (13). The expectation of $\overline{\Xi}^{(\otimes K)}$ is given by
$\mathbb{E}\left[\overline{\Xi}^{(\otimes K)}(\Gamma^{(\otimes K)})\right]$ $=1-\sum_{\mu}\prod_{k=1}^{K}P(X^{(k)}(\tau)=B_{\mu}^{(k)})$
$=1-\sum_{\mu}p_{\mu}(\tau\mid\mu_{0})^{K}.$ (88)
Substituting Eq. (88) into Eq. (17), we obtain
$\sum_{\mu}p_{\mu}(\tau\mid\mu_{0})^{K}\geq e^{-K\tau\mathfrak{a}(t=0)},$ (89)
@@
Appendix D Derivation of Eq. (52)
We show the derivation of Eq. (52), which is the quantum generalization of Eq. (17). Similarly to the classical case, let $\mathfrak{p}_{Q}(\tau)$ be the probability that there is no jump within the interval $0,\tau]$ in $\Gamma_{Q}$. From the result in Ref. [45], the following relation holds:
$\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2}\leq\mathfrak{p}_{Q}(\tau),$ (90)
@@
Here, $\mathcal{B}(\tau)$ is the quantum dynamical activity, whose expression is expressed in Appendix F. We consider the $K$ replica scenario, where there are $K$ identical processes. Let $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ denote the probability that there is no jump in the $K$ replica processes. Since each process is independent, from Eq. (90), $\mathfrak{p}^{(\otimes K)}_{Q}(\tau)$ satisfies
$\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K}\leq\mathfrak{p}_{Q}^{(\otimes K)}(\tau).$ (92)
If no jump occurs in the interval $0,\tau]$, then the observable satisfies $N_{Q\circ}^{(\otimes K)}(\Gamma_{Q\varnothing}^{(\otimes K)})=0$ by the condition in Eq. (47). The converse, however, does not necessarily hold. Therefore, the probability that $N_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{(\otimes K)})=0$ is bounded from below by
$P\left[N_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{(\otimes K)})=0\right]\geq\mathfrak{p}_{Q}^{(\otimes K)}(\tau).$ (93)
@@
$\frac{\mathbb{E}\left[|N_{Q\circ}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}{\mathbb{E}\left[|N_{Q\circ}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}$ $\leq P(|N_{Q\circ}^{(\otimes K)}|>0)$
$=1-P(|N_{Q\circ}^{(\otimes K)}|=0)$
$\leq 1-\mathfrak{p}_{Q}^{(\otimes K)}(\tau)$
@@
where we used Eqs. (92) and (93). Equation (94) proves Eq. (52) in the main text.
Appendix E Derivation of Eq. (57)
We show the derivation of Eq. (57), which directly parallels the derivation of Eq. (27). Specifically, let us consider the following observable for $K$ replicas:
$\overline{\Lambda}_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{\otimes K})\equiv\begin{cases}0&N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\circ}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})\\ 1&\mathrm{otherwise}\end{cases}.$ (95)
Then $\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ returns 0 only when the values of the observable $N_{Q\circ}$ in $K$ processes agree, $N_{Q\circ}^{(1)}(\Gamma_{Q}^{(1)})=N_{Q\mathrm{\circ}}^{(2)}(\Gamma_{Q}^{(2)})=\cdots=N_{Q\circ}^{(K)}(\Gamma_{Q}^{(K)})$. Then the observable $\overline{\Lambda}_{Q\circ}^{(\otimes K)}$ satisfies
$\overline{\Lambda}_{Q\circ}^{(\otimes K)}(\Gamma_{Q\varnothing}^{\otimes K})=0,$ (96)
which is the condition of Eq. (47). Let $\mathcal{Z}_{Q\circ}$ be the set of all possible values that the observable $N_{Q\circ}(\Gamma_{Q})$ can take. Then, the expectation becomes the following:
$\mathbb{E}\left[\overline{\Lambda}_{Q\circ}^{(\otimes K)}(\Gamma_{Q}^{\otimes K})\right]$ $=1-\sum_{z\in\mathcal{Z}_{Q\circ}}\prod_{k=1}^{K}P(N_{Q\circ}^{(k)}(\Gamma_{Q}^{(k)})=z)$
$=1-\sum_{z\in\mathcal{Z}_{Q\circ}}P(N_{Q\circ}(\Gamma_{Q})=z)^{K}.$ (97)
By considering a general $K$ case and substituting Eq. (97) into Eq. (52), we obtain
$\sum_{z\in\mathcal{Z}_{Q\circ}}P(N_{Q\circ}(\Gamma_{Q})=z)^{K}\geq\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K},$ (98)
@@
Appendix F Quantum dynamical activity
We briefly review the quantum dynamical activity, a quantum generalization of the dynamical activity.
@@ Appendix F Quantum dynamical activity
$\mathcal{A}_{Q}(\tau)\equiv\int_{0}^{\tau}\sum_{m}\mathrm{Tr}\left[L_{m}\rho(t)L_{m}^{\dagger}\right]dt.$ (99)
In the classical limit, this expression reduces to the classical dynamical activity given in Eq. (15).
@@
$\mathcal{B}(\tau)=\mathcal{A}_{Q}(\tau)+\mathcal{C}(\tau),$ (100)
@@
Here, $\check{H}(t)\equiv e^{\mathcal{L}^{\dagger}t}H$ is the Hamiltonian $H$ in the Heisenberg picture, $H_{\mathrm{eff}}=H-\frac{i}{2}\sum_{m}L_{m}^{\dagger}L_{m}$ is the effective Hamiltonian, and $\mathcal{L}^{\dagger}$ is the adjoint Liouvillian,